OpenStudy (anonymous):

1/6 x + 2 = 8 for x.

OpenStudy (anonymous):

Same, u isolate x 1/6x+2=8 -2 on both sides 1/6x=6 now flip the fracrion and multiply it on both sides x=36/1 x=36 understand?

OpenStudy (anonymous):

36